Hello,
I do not know how to find the intersection of two lines. Only information I have are the coordinates of those two lines.
Line 1: (0,952)(1111,0)
Line 2: (0,1187)(863,0)
Someone help!

Answer by anand429(138)

I hope you are familiar with slope-intercept form of line
i.e y=mx+c
so for Line 1,
952=0+c (for 1st end point)
=> c=952
and 0=m*1111+c )for 2nd end point)
=> 0=1111*m + 952
=> m= -952/1111
so the eqn. of 1st line is y=(-952/1111)x+952 -(i)
Similarly, eqn of 2nd line is
y=(-1187/863)x + 1187 -(ii)
Solving eqns. (i) and (ii) will give point of intersection.